OpenStudy (anonymous):

the perimeter of a television screen is 112 inches the ratio of the length to the width is 9 to 5 find the length and width length = width=

OpenStudy (radar):

let 9x = length let 5x = width perimeter =18x +10x =28x =112 in. x=4 9x=36 in. length 5x = 20 in width.

OpenStudy (anonymous):

112/2=56inches, this is the length and width add up together. 9+5=14, to see how many ratios are there. 56/14=4, this is one ratio. length is 5*4=20inches, width is 4*4=16inches

OpenStudy (anonymous):

thank you very much :)

OpenStudy (radar):

@e09876w with a width of 16 and a length of 20 is the perimeter 112 in.?

OpenStudy (radar):

40 + 32 is 72 in. not 112 in.

OpenStudy (radar):

nisa did you follow with understanding my solution?

OpenStudy (radar):

Proof does the perimeter = 112 in. 36 + 36 +20 +20 =112 112=112 Checks/ Is there a ratio of 9 to 5? 36/20=9/5 checks. The answer is correct.

OpenStudy (anonymous):

yes i did get it thankz alot i appreciate it

OpenStudy (radar):

ur welcome and good luck with your studies.

OpenStudy (anonymous):

:) good luck with whatever your doing in life :)